56) Social science. In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were cast in all, how many votes did each candidate receive?



Answer by checkley75(3666)

X+(X+220)=810
X+X+220=810
2X=810-220
2X=590
X=590/2
X=295 VOTES FOR THE LOOSER.
295+220=515 VOTES FOR THE WINNER.
PROOF
295+515=810
810=810